# NuLite - Lightweight and Fast Model for Nuclei Instance Segmentation and Classification

### v1

@@ 4.5 Models complexity analysis
To prove that our model, NuLite, has a lower complexity than CellViT, in Table 6, we report an exhaustive comparison between NuLite and CellViT in terms of parameters count and GFlops, estimated size, and latency on GPU using an input shape of 256 and 1024. In particular, we consider all FastViT models (T8, T12, S12, SA24, SA36, and MA36) and the reparameterized versions. Instead, for CellViT, we use the version with ViT256, SAM-B, SAM-L, and SAM-H as encoders. Results concerning GFlops and estimated size refer to a batch with just one image. Instead, GPU latency refers to a batch size of 4, and we repeated the experiments 100 times and reported the mean and variance in milliseconds. We conducted the measure on a server with AMD EPYC 7282 16-Core Processor, RAM 64 GB, and GPU Nvidia Tesla V100S 32 GB. We consider reparameterization because even if the number of parameters and GFlops are roughly the same, the inference time is lower using it in the inference step, so we limit our consideration of these results in this analysis. According to these results, we can note that in terms of GFLOPS, all versions of NuLite are significantly lower than CellViT; further, all NuLite sizes are lower than CellViT. In terms of parameters, CellViT with SAM as the backbone is really big compared with our model, but the version with ViT-256 is smaller than NuLite with FastViT-MA36 as the backbone. CellViT takes a longer GPU latency than our NuLite versions because the amount of multiplication and addition is smaller than all CellViT versions. Moreover, if we consider our worst NuLite GPU latency for each shape, namely using FastViT-MA36 without reparameterization, it is faster than the best case of CellViT, namely CellViT256; moreover, it is almost two times faster with shape 1024x1024. Limiting the analysis only to selected NuLite, we compared than with CellViT256 and CellViT-SAM-H in terms of the parameters, GFLOPS, and GPU latency. Table 7 presents a comparative analysis of the GFLOPS and the number of parameters for the NuLite models against CellViT architectures, including CellViT-256 and CellViT-SAM-H. The values indicate how computationally intensive and parameter-rich the CellViT models are compared to the corresponding NuLite variants. For GFLOPS, the CellViT-256 model is approximately 4.4 to 5.1 times more intensive than NuLite models, while CellViT-SAM-H ranges from 7.2 to 8.2 times more. Regarding the number of parameters, CellViT-256 has 1.1 to 2.7 times more parameters, whereas CellViT-SAM-H exhibits a significantly larger increase, ranging from 17 to 40.8 times more than NuLite models.
Table 7: Comparison of GFLOPS and parameters for NuLite models highlighting the multiples of GFLOPS and the number of parameters across CellViT architectures.
GFLOPS # Parameters
@@
NuLite-T 5.1×5.1$\times$5.1 × 8.2×8.2$\times$8.2 × 2.7×2.7$\times$2.7 × 40.8×40.8$\times$40.8 ×
NuLite-M 4.7×4.7$\times$4.7 × 7.6×7.6$\times$7.6 × $\times2$ $\times3$
NuLite-H $\times4$ 7.2×7.2$\times$7.2 × $\times1$ 17×17$\times$17 ×
Table 8 compares the inference speedups of NuLite models relative to CellViT architectures, specifically CellViT-256 and CellViT-SAM-H, for different patch sizes (256x256 and 1024x1024 pixels). The values indicate how much faster the NuLite models perform than the CellViT models. For a 256x256 patch size, the NuLite-T model achieves a speedup of 2.7 times over CellViT-256 and 7.8 times over CellViT-SAM-H. The NuLite-M and NuLite-H models also demonstrate notable speed improvements, though to a lesser extent. For the larger patch size of 1024x1024 pixels, the speedup is even more pronounced; NuLite-T reaches a speedup of 4.6 times over CellViT-256 and 7.7 times over CellViT-SAM-H.
Table 8: Inference Speedups for NuLite models with patch size 256X256 pixel and overlap 64px and patch size 1024X1024 over and CellViT architectures.
Patch size: 256 X 256 Patch size: 1024 X 1024
@@
NuLite-T 2.7×2.7$\times$2.7 × 7.8×7.8$\times$7.8 × 4.6×4.6$\times$4.6 × $\times7$
NuLite-M 2×2$\times$2 × 5.9×5.9$\times$5.9 × 4×4$\times$4 × $\times6$
NuLite-H 1.6×1.6$\times$1.6 × 4.7×4.7$\times$4.7 × 3.5×3.5$\times$3.5 × 5.9×5.9$\times$5.9 ×
Lastly, to prove that our models are lighter than CellViT but performing as well as it, Figure 4 shows a comparison between them in terms of mPQ and bPQ, each error bar is the average standard deviation over tissue, on the x-axis there are GPLOS on log-scale. Analyzing this image, we can note that our models are less complex than CellViT variants, but the results are approximately the same.
@@
This indicates that while NuLite models maintain a lower computational and parameter footprint, they are being compared to highly demanding SOTA architectures, emphasizing their efficiency. Furthermore, the consistent performance advantage highlights the efficiency of NuLite models in inference speed, particularly when dealing with larger image patches.
4.6 Results on others datasets
Figure 5: Segmentation masks generated by CellViT256, NuLite-T, NuLite-H, CellViT-SAM-H, and NuLite-M models on a histological image of tissue from the MoNuSeg dataset. Models were evaluated at 256x256 and 1024x1024 resolutions. Masks highlight nuclei.
Table 9: Comparison of CellViT, CellViT-SAM-H, and NuLite models (NuLite-T, NuLite-M, NuLite-H) across MoNuSeg, CoNSeP, and GlySAC datasets. Metrics include Detection Quality (DQ), Segmentation Quality (SQ), Panoptic Quality (PQ), and detection precision ($P_{d}$), recall ($R_{d}$), and F1-score ($F_{1,d}$) for patch sizes of 256x256 px and 1024x1024 px. The best results are highlighted in bold, with the second-best in underlined text.
Model Patch-Size: 256 x 256 px - Overlap: 64 px Patch-Size: 1024x1024 px
@@ 4.6 Results on others datasets
To understand the capability of generalization of NuLite, we used MoNuSeg, CoNSeP, and GlySAC datasets and compared the results against CellViT. In particular, we used GlySAC and CoNSeP to evaluate segmentation and classification performance. Instead, we used MoNuSeg to evaluate only segmentation performance because it does not provide nuclei type. As described in the dataset section, CoNSeP and GlySAC have different nuclei types of PanNuke, so we aligned them to compute multi-class metrics. All datasets contain tiles with shape 1000x1000, following the workflow introduced in [11], we resized them to 1024x1024 pixels. Lastly, we compared the results using the input shape of 256x256 pixels with an overlap of 64 pixels and 1024x1024 pixels. The authors in [11] proved that using an input shape of 1024x1024 does not negatively affect the results, but they analyzed what changes for multi-class metrics; in this section, we also analyze this aspect to understand if it is possible to use 1024x1024 pixels tile as input when we use NuLite on whole slide images. Here, we first analyze the binary metrics; table 9 contains the Detection Quality (DQ), Segmentation Quality (SQ), Panoptic Quality (PQ), Precision ($P_{d}$), Recall ($R_{d}$), and F1-score ($F_{1,d}$) for each dataset and inference configuration. First, we can confirm that using a tile of 1024x1024 as input is roughly equivalent to using a tile of 256x256 pixels with an overlap of 64 pixels.
NuLite consistently demonstrates competitive or superior performance compared to CellViT, particularly in precision and recall metrics. For instance, in the MoNuSeg dataset with 256x256 patches, NuLite-T achieves the highest recall ($R_{d}$) of 0.910, suggesting a strong ability to identify relevant instances. Similarly, NuLite-H delivers the best recall (0.725) and F1 score (0.774) for the CoNSeP dataset with smaller patches. Although CellViT-SAM-H shows slightly better overall scores on some datasets, NuLite variants are often close. NuLite-M even outperforms certain metrics, such as precision ($P_{d}$) in the GlySAC dataset. These findings suggest that NuLite, especially its medium and high variants, can match or surpass the performance of the current state-of-the-art model, CellViT, in specific scenarios. Concerning binary results, we also show a visual example in Figure 5; in particular, it contains a tile of an image of the MoNuSeg date and an inference example of each analyzed model and each inference configuration setting.
Figure 6: Segmentation masks generated by CellViT256, NuLite-T, NuLite-H, CellViT-SAM-H, and NuLite-M models on a histological image of epithelial tissue from the CoNSeP dataset. Models were evaluated at 256x256 and 1024x1024 resolutions. Masks highlight neoplastic, inflammatory, epithelial, and miscellaneous regions.
Table 10: Performance of CellViT-256, CellViT-SAM-H, NuLite-T, NuLite-M, and NuLite-H on the CoNSeP dataset across Neoplastic, Epithelial, Inflammatory, and Miscellaneous nuclei type with two patch sizes (256x256 px and 1024x1024 px). Metrics include Detection Quality (DQ), Segmentation Quality (SQ), and Panoptic Quality (PQ), along with detection precision ($P_{d}$), recall ($R_{d}$), and F1-score ($F_{1,d}$). The best results are highlighted in bold, and the second-best results are underlined.
Model Patch-Size: 256 x 256 px - Overlap: 64 Patch-Size: 1024 x 1024 px
@@
[…]
Concerning the CoNSeP multi-class setting, we followed the alignment as shown in [29]; the Neoplastic class includes PanNuke’s neoplastic and CoNSeP’s dysplastic/malignant epithelial; the Inflammatory class encompasses PanNuke’s inflammatory and CoNSeP’s inflammatory; the Epithelial class consists of PanNuke’s epithelial and CoNSeP’s healthy epithelial; finally, the Miscellaneous class incorporates PanNuke’s dead and connective tissues alongside CoNSeP’s other types, which include fibroblast, muscle, and endothelial tissues.
Table 10 reports the results for CoNSeP multi-class comparing CellViT variants and the proposed NuLite variants across multiple tissue classes: Neoplastic, Epithelial, Inflammatory, and Miscellaneous. Each model performance is evaluated using the two configuration settings described above, with metrics such as DQ (Dice Quality), SQ (Segmentation Quality), PQ (Panoptic Quality), Precision ($P_{d}$), Recall ($R_{d}$), and F1-score ($F_{1,d}$) reported for each class.
For patch size 256x256 pixels with 64-pixel overlap, the NuLite-H model demonstrates competitive performance, tying for the top score in the Neoplastic category (0.55) and performing strongly across other categories. CellViT-SAM-H generally leads in this setting, indicating its effectiveness with smaller image patches. However, the NuLite models show notable strengths in specific categories, highlighting their robustness and versatility. For patch size 1024x1024 pixels, NuLite models, particularly NuLite-H, show significant improvement and often outperform CellViT models. Notably, NuLite-H achieves the highest PQ scores in the Neoplastic category (0.58) and ties for the highest in the Miscellaneous category (0.44).
Regarding $F_{1,d}$ scores, NuLite-H consistently excels, particularly with larger patch sizes achieving top scores in the Neoplastic (0.62), Epithelial (0.67), and Miscellaneous (0.59) nuclei type. This indicates that NuLite-H captures high precision and maintains strong recall, which is essential for accurate and reliable classification. The performance of NuLite-H in larger patches is particularly noteworthy, as it highlights the ability of the model to integrate broader contextual information, which is critical in complex image analysis tasks such as histopathology.
Furthermore, Figure 6 shows an inference example for each analyzed model highlighting a tile of an image of the CoNSeP dataset, where we can note that NuLite-H outputs the best visual result.
@@
Table 11: Performance metrics for CellViT-256, CellViT-SAM-H, NuLite-T, NuLite-M, and NuLite-H on the GlySAC dataset. Metrics are provided for different patch sizes (256x256 px and 1024x1024 px) and include Detection Quality (DQ), Segmentation Quality (SQ), and Panoptic Quality (PQ) for Epithelial, Inflammatory, and Miscellaneous categories. Detection precision ($P_{d}$), recall ($R_{d}$), and F1-score ($F_{1,d}$) are also shown. The highest values are highlighted in bold, and the second-highest values are underlined.
Model Patch-Size: 256 x 256 px - Overlap: 64 Patch-Size: 1024 x 1024 px
@@
[…]
Concerning the results on GlySAC, we aligned the nuclei types as follows: the Epithelial class of GlySAC with Neoplatist and Epithelial of PanNuke, the Inflammatory class perfectly match, and the other class of PanNuke with the miscellaneous class of GlySAC. The results in Table 11 highlight the performance differences between NuLite variants and CellViT. The key metrics evaluated are Detection Quality (DQ), Segmentation Quality (SQ), Panoptic Quality (PQ), Precision, Recall, and F1 score for each nucleus type.
@@
For epithelial nuclei, NuLite-H and CellViT-SAM-H exhibit similar high performance with Panoptic Quality (PQ) values of 0.42-0.43 at the 256x256 px patch size. At the larger 1024x1024 px input size, NuLite-H slightly outperforms with a PQ of 0.42, while NuLite-M and CellViT models are closely aligned. Regarding inflammatory nuclei, NuLite-M achieves the highest PQ score of 0.40 for 1024x1024 px patches, indicating superior performance. NuLite-H and CellViT-SAM-H also show strong results with PQ values around 0.39-0.41, demonstrating that NuLite effectively handles inflammatory cell segmentation, though the differences with CellViT are minimal. For miscellaneous nuclei, NuLite-M stands out with the highest PQ of 0.24 at the 1024x1024 px patch size, suggesting better segmentation performance in this challenging category. The advantage is less pronounced at the 256x256 px patch size, but NuLite still performs competitively. Furthermore, NuLite-H demonstrates strong performance in F1 scores, particularly in the 1024x1024 px input size, where it achieves the highest scores across most categories. This model reaches an F1 score of 0.50 for epithelial cells, indicating an effective balance between precision and recall. NuLite-T achieves a high recall value of 0.58 for epithelial cells with the 256x256 px patch size, indicating its effectiveness in identifying a larger proportion of relevant instances. However, this increased sensitivity may come at the cost of precision. Lastly, Figure 7 shows an inference example on GlySAC, where we can observe that NuLite-H achieves good results compared to ground truth.

### v2

@@ 4.5 Models complexity analysis
To prove that our model, NuLite, has a lower complexity than CellViT, in Table 6, we report an exhaustive comparison between NuLite and CellViT in terms of parameters count and GFlops, estimated size, and latency on GPU using an input shape of 256 and 1024. In particular, we consider all FastViT models (T8, T12, S12, SA24, SA36, and MA36) and the reparameterized versions in the inference step. Instead, for CellViT, we use the version with ViT256 and SAM-H as encoders. Results concerning GFlops and estimated size refer to a batch with just one image. Instead, GPU latency refers to a batch size of 4, and we repeated the experiments 100 times and reported the mean and variance in milliseconds. We conducted the measure on a server with AMD EPYC 7282 16-Core Processor, RAM 64 GB, and GPU Nvidia Tesla V100S 32 GB. We consider reparameterization because even if the number of parameters and GFlops are roughly the same, the inference time is lower using it in the inference step, so we limit our consideration of these results in this analysis. According to these results, in terms of GFLOPS, all versions of NuLite are significantly lower than CellViT; further, all NuLite sizes are lower than CellViT. In terms of parameters, CellViT with SAM as the backbone is larger than our model, but the version with ViT-256 is smaller than NuLite with FastViT-MA36 as the backbone. CellViT takes a longer GPU latency than our NuLite versions because the amount of multiplication and addition is smaller than all CellViT versions. Moreover, if we consider our worst NuLite GPU latency for each shape, namely using FastViT-MA36 without reparameterization, it is faster than the best case of CellViT, namely CellViT256; moreover, it is almost two times faster with shape $1024\times 1024$. Furthermore, another critical aspect is the estimated total size, which indicates the amount of memory used by the model during an inference step on a batch size of one, where our NuLite models outperform all CellViT models. Limiting the analysis only to selected NuLite, we compared them with CellViT256 and CellViT-SAM-H in terms of the parameters, GFLOPS, GPU latency, and estimated size.
Table 7: Speedup of NuLite compared to CellViT
(a) #Parameters
CellViT-256 CellViT-SAM-H
NuLite-T 3.89×3.89$\times$3.89 × 58.27×58.27$\times$58.27 ×
NuLite-M 1.37×1.37$\times$1.37 × 20.56×20.56$\times$20.56 ×
NuLite-H 0.98×0.98$\times$0.98 × 14.62×14.62$\times$14.62 ×
(b) GFLOPS
CellViT-256 CellViT-SAM-H
NuLite-T 6.73×6.73$\times$6.73 × 10.84×10.84$\times$10.84 ×
NuLite-M 5.74×5.74$\times$5.74 × 9.26×9.26$\times$9.26 ×
NuLite-H 4.08×4.08$\times$4.08 × 6.58×6.58$\times$6.58 ×
Table 7 presents a comparative analysis of the number of parameters for the NuLite models against CellViT architectures, including CellViT-256 and CellViT-SAM-H. The values indicate how parameter-rich the CellViT models are to the corresponding NuLite variants. CellViT-256 has 3.89 to 0.89 times the number of parameters of NuLite, whereas CellViT-SAM-H exhibits a significantly larger increase, from 14.62 to 58.27 times more than NuLite models. In the same way, Table 7(b) presents the comparative analysis for GFLOPS; the CellViT-256 model is 4.08 to 6.73 times more intensive than NuLite models, while CellViT-SAM-H ranges from 6.58 to 10.84 times more.
Table 8: Estimated size speedups for NuLite models with input size $256\times 256$ pixels with overlap 64 pixels and input size $1024\times 1024$ pixel against CellViT models.
Input Size $256\times 256$ Input Size $1024\times 1024$
@@
NuLite-T 3.8×3.8$\times$3.8 × 12.25×12.25$\times$12.25 × 3.79×3.79$\times$3.79 × 6.41×6.41$\times$6.41 ×
NuLite-M 2.49×2.49$\times$2.49 × 8.05×8.05$\times$8.05 × 2.73×2.73$\times$2.73 × 4.61×4.61$\times$4.61 ×
NuLite-H 2.04×2.04$\times$2.04 × 6.57×6.57$\times$6.57 × 2.29×2.29$\times$2.29 × 3.88×3.$8$
Table 9: Inference speedups for NuLite models with input size $256\times 256$ pixels with overlap 64 pixels and input size $1024\times 1024$ pixel against CellViT models.
Input Size $256\times 256$ Input Size $1024\times 1024$
@@
NuLite-T 3.58×3.58$\times$3.58 × 10.43×10.43$\times$10.43 × 6.19×6.19$\times$6.19 × 12.64×12.64$\times$12.64 ×
NuLite-M 1.91×1.91$\times$1.91 × 5.57×5.57$\times$5.57 × 4.38×4.38$\times$4.38 × 8.95×8.95$\times$8.95 ×
NuLite-H 1.55×1.$5$ 4.51×4.51$\times$4.51 × 2.9×2.9$\times$2.9 × 5.93×5.93$\times$5.93 ×
@@
Again, Table 8 presents the comparative analysis for estimated size during the inference step; for input size $256\times 256$, the CellViT-256 model consumes from 2.04 to 3.8 times more memory than NuLite models, while CellViT-SAM-H ranges from 6.57 to 12.25 times more; for input size $1024\times 1024$, the CellViT-256 model consumes from 2.29 to 3.79 times more memory than NuLite models, while CellViT-SAM-H ranges from 3.88 to 6.41 times more.
Figure 5: Segmentation masks generated by CellViT256, NuLite-T, NuLite-H, CellViT-SAM-H, and NuLite-M models on a histological image of tissue from the MoNuSeg dataset. Models were evaluated at $256\times 256$ and $1024\times 1024$ resolutions. Masks highlight nuclei.
Table 10: Comparison of CellViT, CellViT-SAM-H, and NuLite models (NuLite-T, NuLite-M, NuLite-H) across MoNuSeg, CoNSeP, and GlySAC datasets. Metrics include Detection Quality (DQ), Segmentation Quality (SQ), Panoptic Quality (PQ), and detection precision ($P_{d}$), recall ($R_{d}$), and F1-score ($F_{1,d}$) for patch sizes of $256\times 256$ px and $1024\times 1024$ px. The best results are highlighted in bold, with the second-best in underlined text.
Model Patch-Size: $256\times 256$ px - Overlap: 64 px Patch-Size: $1024\times 1024$ px
@@
[…]
Table 9 compares the inference speedups of NuLite models relative to CellViT architectures, specifically CellViT-256 and CellViT-SAM-H, for different patch sizes ($256\times 256$ and $1024\times 1024$ pixels). The values indicate how much faster the NuLite models perform than the CellViT models. For a $256\times 256$ patch size, the NuLite models speed up from 1.55 to 3.58 times CellViT-256 and from 4.51 to 10.43 times over CellViT-SAM-H. For a $1024\times 1024$ patch size, the NuLite models speed up from 2.9 to 6.19 times CellViT-256 and from 5.93 to 12.64 times over CellViT-SAM-H. Lastly, to prove that our models are lighter than CellViT but performing as well as it, Figure 4 shows a comparison between them in terms of mPQ and bPQ, each error bar is the average standard deviation over tissue, on the x-axis there are GPLOS on log-scale. Analyzing this image, we can note that our models are less complex than CellViT variants, but the results are approximately the same. These aspects indicate that NuLite models maintain a lower computational and parameter footprint than highly demanding SOTA architectures, emphasizing their efficiency. Furthermore, the consistent performance advantage highlights the efficiency of NuLite models in inference speed, mainly when dealing with larger image patches.
Figure 6: Segmentation masks generated by CellViT256, NuLite-T, NuLite-H, CellViT-SAM-H, and NuLite-M models on a histological image of epithelial tissue from the CoNSeP dataset. Models were evaluated at $256\times 256$ and $1024\times 1024$ resolutions. Masks highlight neoplastic, inflammatory, epithelial, and miscellaneous regions.
Table 11: Performance of CellViT-256, CellViT-SAM-H, NuLite-T, NuLite-M, and NuLite-H on the CoNSeP dataset across Neoplastic, Epithelial, Inflammatory, and Miscellaneous nuclei type with two patch sizes ($256\times 256$ px and $1024\times 1024$ px). Metrics include Detection Quality (DQ), Segmentation Quality (SQ), and Panoptic Quality (PQ), along with detection precision ($P_{d}$), recall ($R_{d}$), and F1-score ($F_{1,d}$). The best results are highlighted in bold, and the second-best results are underlined.
Model Patch-Size: $256\times 256$ px - Overlap: 64 Patch-Size: 1024 x 1024 px
@@
4.6 Results on others datasets
To understand the capability of generalization of NuLite, we used MoNuSeg, CoNSeP, and GlySAC datasets and compared the results against CellViT. In particular, we used GlySAC and CoNSeP to evaluate segmentation and classification performance. Instead, we used MoNuSeg to evaluate only segmentation performance because it does not provide nuclei type. As described in the dataset section, CoNSeP, and GlySAC have different nuclei types of PanNuke, so we aligned them to compute multi-class metrics. All datasets contain tiles with shape 1000x1000, following the workflow introduced in [12], we resized them to $1024\times 1024$ pixels. Lastly, we compared the results using the input shape of $256\times 256$ pixels with an overlap of 64 pixels and $1024\times 1024$ pixels. The authors in [12] proved that using an input shape of $1024\times 1024$ does not negatively affect the results, but they analyzed what changes for multi-class metrics; in this section, we also analyze this aspect to understand if it is possible to use $1024\times 1024$ pixels tile as input when we use NuLite on whole slide images. Here, we first analyze the binary metrics; table 10 contains the Detection Quality (DQ), Segmentation Quality (SQ), Panoptic Quality (PQ), Precision ($P_{d}$), Recall ($R_{d}$), and F1-score ($F_{1,d}$) for each dataset and inference configuration. First, we can confirm that using a tile of $1024\times 1024$ as input is roughly equivalent to using a tile of $256\times 256$ pixels with an overlap of 64 pixels.
For the MoNuSeg Dataset with a smaller patch size (256x256), although CellViT-SAM-H demonstrates the best overall performance, leading in DQ, SQ, PQ, and precision, our solution, NuLite-T, achieves the highest recall and F1-score, highlighting its strong detection capabilities. NuLite-H also exhibits competitive performance, closely following CellViT-SAM-H in several key metrics. When considering the larger patch size (1024x1024), while CellViT-SAM-H continues to outperform others, NuLite, particularly NuLite-T, remains highly competitive, performing closely across multiple evaluation metrics.
On the CoNSeP Dataset, with a 256x256 patch size, CellViT-SAM-H again leads in DQ, SQ, and PQ and shows the best recall. However, NuLite-M excels in precision and F1-score, underlining its superior detection accuracy. With the larger patch size of 1024x1024, CellViT-SAM-H continues to dominate most categories, especially in DQ, PQ, recall, and F1-scores. Still, NuLite-M achieves the highest precision and maintains robust performance across other metrics.
For the GlySAC Dataset with a 256x256 patch size, CellViT-256 slightly outperforms others in DQ, SQ, and PQ. Nevertheless, our NuLite-M demonstrates strong precision and F1-score, leading in precision. At the larger patch size of 1024x1024, CellViT-256 and NuLite-T exhibit the best performance in DQ and PQ, with CellViT-256 achieving the highest scores overall, while NuLite-M outperforms others in both F1-score and precision. Concerning binary results, we also show a visual example in Figure 5; in particular, it contains a tile of an image of the MoNuSeg date and an inference example of each analyzed model and each inference configuration setting.
@@ 4.6 Results on others datasets
Table 12: Performance metrics for CellViT-256, CellViT-SAM-H, NuLite-T, NuLite-M, and NuLite-H on the GlySAC dataset. Metrics are provided for different patch sizes ($256\times 256$ px and $1024\times 1024$ px) and include Detection Quality (DQ), Segmentation Quality (SQ), and Panoptic Quality (PQ) for Epithelial, Inflammatory, and Miscellaneous categories. Detection precision ($P_{d}$), recall ($R_{d}$), and F1-score ($F_{1,d}$) are also shown. The highest values are highlighted in bold, and the second-highest values are underlined.
Model Patch-Size: $256\times 256$ px - Overlap: 64 Patch-Size: $1024\times 1024$ px
@@
NuLite-H 0.465 0.534 0.462 0.513 0.452 0.418 0.318 0.309 0.278 0.468 0.525 0.451 0.523 0.441 0.417 0.328 0.326 0.291
Concerning the CoNSeP multi-class setting, we followed the alignment as shown in [30]; the Neoplastic class includes PanNuke’s neoplastic and CoNSeP’s dysplastic/malignant epithelial; the Inflammatory class encompasses PanNuke’s inflammatory and CoNSeP’s inflammatory; the Epithelial class consists of PanNuke’s epithelial and CoNSeP’s healthy epithelial; finally, the Miscellaneous class incorporates PanNuke’s dead and connective tissues alongside CoNSeP’s other types, which include fibroblast, muscle, and endothelial tissues. Table 11 reports the results for CoNSeP multi-class comparing CellViT variants and the proposed NuLite variants across multiple tissue classes: Neoplastic, Epithelial, Inflammatory, and Miscellaneous. Each model performance is evaluated using the two configuration settings described above, with metrics such as Detection Quality (DQ), Segmentation Quality (SQ), Panoptic Quality (PQ), Precision ($P_{d}$), Recall ($R_{d}$), and F1-score ($F_{1,d}$) reported for each class. For patch size $256\times 256$ pixels with a 64-pixel overlap, the NuLite-H model shows competitive performance, tying for the top score in the Neoplastic category for both DQ (0.571) and PQ (0.442) and demonstrating strong results across other categories. CellViT-SAM-H generally leads in this setting, indicating its effectiveness with smaller image patches, particularly in the Epithelial category, where it achieves the highest DQ (0.772), SQ (0.79), and PQ (0.61) scores. However, the NuLite models, particularly NuLite-M and NuLite-H, show notable strengths in specific categories, highlighting their robustness and versatility. When it comes to larger patch sizes, particularly $1024\times 1024$ pixels, the NuLite models, especially NuLite-H, show a significant improvement and often outperform CellViT models. NuLite-H, in particular, achieves the highest PQ scores in the Neoplastic category (0.456) and ties for the highest in the Epithelial category (0.601), demonstrating its ability to maintain performance across larger contexts. NuLite-M also performs exceptionally well, achieving the highest PQ score in the Miscellaneous category (0.438) and the top DQ score in the Inflammatory category (0.665). Regarding $F_{1,d}$ scores, NuLite-M consistently excels, particularly with larger patch sizes, achieving top scores in the Neoplastic (0.664), Epithelial (0.667), and Miscellaneous (0.605) nuclei types. This indicates that NuLite-M captures high precision while maintaining strong recall, essential for accurate and reliable classification. Furthermore, Figure 6 shows an inference example for each analyzed model, highlighting a tile of an image from the CoNSeP dataset. The visual results indicate that NuLite-H and NuLite-M output the best segmentation masks, particularly in more challenging regions, underscoring their effectiveness in histological image analysis.
@@
Concerning the results on GlySAC, we aligned the nuclei types as follows: the Epithelial class of GlySAC with Neoplatist and Epithelial of PanNuke, the Inflammatory class perfectly match, and the other class of PanNuke with the miscellaneous class of GlySAC. The results in Table 12 highlight the performance differences between NuLite variants and CellViT. The key metrics evaluated are Detection Quality (DQ), Segmentation Quality (SQ), Panoptic Quality (PQ), Precision, Recall, and F1 score for each nucleus type. For epithelial nuclei, the NuLite-M model leads with the highest PQ score of 0.431 at the $256\times 256$ px patch size. At the $1024\times 1024$ px size, NuLite-M continues to excel with the highest PQ of 0.429, slightly outperforming both CellViT-SAM-H and NuLite-H, which also deliver strong results. Regarding inflammatory nuclei, CellViT-SAM-H outperforms all other models at the $256\times 256$ px size, achieving the highest PQ of 0.415. However, at the $1024\times 1024$ px patch size, NuLite-M takes the lead with a PQ of 0.395, slightly ahead of CellViT-SAM-H and NuLite-H. In the miscellaneous nuclei category, NuLite-M again stands out, particularly at the $1024\times 1024$ px patch size, where it achieves the highest PQ score of 0.222. This suggests that NuLite-M handles the challenging task of segmenting miscellaneous nuclei more effectively than the other models, especially with larger patches. Examining the F1-scores across the models, NuLite-M consistently demonstrates strong performance, particularly in the $1024\times 1024$ px patch size, where it achieves the highest F1-scores across most categories. Notably, it reaches an F1-score of 0.501 for epithelial cells, indicating a well-balanced performance between precision and recall. CellViT-SAM-H also shows competitive performance with high F1-scores, particularly for inflammatory nuclei at the smaller patch size. Regarding precision and recall, NuLite-M has the highest precision in the epithelial and miscellaneous categories at the $256\times 256$ px patch size. In contrast, CellViT-SAM-H has the highest recall for inflammatory nuclei. This trend is consistent at the $1024\times 1024$ px patch size, where NuLite-M maintains high precision across most categories. Lastly, Figure 7 shows an inference example on GlySAC, where we can observe that NuLite-H achieves good results compared to ground truth.